# JMLR: Joint Medical LLM and Retrieval Training for Enhancing Reasoning and Professional Question Answering Capability

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated a remarkable potential in medical knowledge acquisition and question-answering. However, LLMs can potentially hallucinate and yield factually incorrect outcomes, even with domain-specific pretraining. Previously, retrieval augmented generation (RAG) has limited success in addressing hallucinations. Unlike previous methods in RAG where the retrieval model was trained separately from the LLM, we introduce JMLR (for Jointly trains LLM and information Retrieval (IR)) during the fine-tuning phase. The synchronized training mechanism enhances JMLR’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions and reduces the demand for computational resources. We evaluated JMLR on the important medical question answering application. Our experimental results demonstrate that JMLR-13B (70.5%) outperforms a previous state-of-the-art open-source model using conventional pre-training and fine-tuning Meditron-70B (68.9%) and Llama2-13B with RAG (54.9%) on a medical question-answering dataset. JMLR-13B (148 GPU hours) also trains much faster than Meditron-70B (42630 GPU hours). Through this work, we provide a new and efficient knowledge enhancement tool for healthcare, demonstrating the potential of integrating IR and LLM training for medical question-answering systems  [^1] The code(https://github.com/believewhat/JMLR-Joint-Medical-LLM-and-Retrieval-Training), along with selected retrieval data that can be made public, is included in the supplementary material and will be made publicly accessible with CC-BY 4.0 license upon the paper’s acceptance..
@@ 1 Introduction
Figure 1: JLMR achieved the highest average accuracy across the MMLU-Medical, MedMcQA, MedQA, and Amboss datasets, utilizing only 148 GPU hours.
@@
Effective clinical decision-making relies on a logical diagnostic chain, which requires specialized knowledge that isn’t widely accessible, presenting a major healthcare challenge. Large Language Models (LLMs) show promise in making this critical medical knowledge more accessible (Peng et al., 2023; Yang et al., 2023). Traditionally, LLMs have been developed for general tasks using data from diverse online sources, leading to a lack of high-quality, domain-specific information, especially in medicine Wornow et al. (2023); Singhal et al. (2022). Previous researches expand medical knowledge by continue training general domain LLMs with domain-specific datasets. (Yunxiang et al., 2023; Zhang et al., 2023; Toma et al., 2023).
However, LLMs face the significant challenge of “hallucination”, where models generate plausible but incorrect or unverified information (Ji et al., 2023; Bang et al., 2023). Such errors raise serious concerns in healthcare, where accuracy is imperative (Ahmad et al., 2023). Hence, recent works proposed retrieval augmented generation (RAG): they first train a retriever to obtain relevant documents from a general domain corpus based on the input query and then train an LLM to generate a response based on the input query and the retrieved documents  (Lewis et al., 2020; Borgeaud et al., 2022a; Cheng et al., 2023; Xiong et al., 2024). By retrieving domain specific documents, RAG identifies relevant knowledge and provides contextual grounding for LLMs, thereby alleviating hallucination issues  (Shuster et al., 2021; Zhu et al., 2023).
Although RAG shows high accuracy in benchmarks for open-domain question answering, its effectiveness in specific domains is yet in challenge because retrievers trained in the general domain usually perform worse than those fine-tuned in the specific domain  (Gao et al., 2023; Zhao et al., 2022; Thakur et al., 2021). Fine-tuning retrievers requires document-query pairs, which may not be readily available for a specific domain. Labeling such medical pairs specifically for this purpose incurs additional time and financial costs. Moreover, RAG trains retriever and LLM separately. When training LLM with a frozen retriever, the retrieval may not be optimized to align with LLM to produce the correct answer (Asai et al., 2024; Yoran et al., 2023; Rubin et al., 2022). In this study, we introduce a novel method to jointly train retrieval and LLM. Although the method is generalizable, we focus on the evaluation on the important medical domain where the accuracy of the generated content is crucial.
@@
To validate the effectiveness of JMLR in terms of reducing computational resource requirements and fully utilizing given knowledge, we designed three experimental setups. These experiments evaluated whether direct fine-tuning, instead of traditional pretraining, would improve JMLR’s performance. We utilized external data resources for retrieval, including the MIMIC-IV dataset (Johnson et al., 2020), medical textbooks, and diverse medical documents for knowledge expansion. We experimented with different LLMs, including PMC-Llama, Meditron, ChatGPT with or without MedRAG. Our JMLR-7B has demonstrated superior results, achieving an impressive 62.45% accuracy, outperforming the traditional method’s 53.2% on the medical QA dataset. Furthermore, our 13B model (69.2%) surpasses the performance of both the open-source medical LLM (Meditron 70B: 68.9%) and closed-source general LLM (ChatGPT: 58.5%). Moreover, evaluation by both GPT-4 and domain experts supports the superior performance of JMLR. We summarize our key contributions as follows:
  • We propose a new method that jointly train retrieval and LLM, resulting in JMLR surpassing the state-of-the-art open source model.
  • By training jointly, we do not require additional human annotation, thereby enhancing flexibility in domain or downstream task adaptation.
@@ 3.1 Retriever
For information retrieval we used the ColBERT model and the trained weight as the initial model. ColBERT utilizes BERT-based encoders to transform both queries and documents into bags of embeddings. A single BERT model is shared between the query and document encoders, but the inputs are distinguished by appending special tokens: $[Q]$ for queries and $[D]$ for documents. The text is tokenized for queries into BERT-based WordPiece tokens $q_{0},q_{1}\ldots q_{l}$, with the $[Q]$ token placed after BERT’s sequence-start token $[CLS]$. For documents, ColBERT segments a document d into its constituent tokens $d_{1},d_{2},\ldots d_{n}$, to which we prepend BERT’s start token $[CLS]$ followed by our special token $[D]$ that indicates a document sequence. Queries shorter than a predefined number of tokens $N_{q}$ are padded with BERT’s special $[mask]$ tokens to reach the length $N_{q}$, otherwise truncated. The padded sequence of input tokens is then passed through BERT. The resulting embeddings for queries ($E_{q}$) and documents ($E_{d}$) are computed as follows: $E_{q}$ is normalized after being processed through a CNN layer applied to BERT’s output, and $E_{d}$ undergoes a similar process but includes a filtering step to remove certain embeddings.
@@ 3.2 LLM
We adapted Llama as our LLM model. The original Llama model supports only a maximum of 4k tokens, which poses a challenge when multiple documents are retrieved. To address this, we utilized the Shifted Sparse Attention (S2-Attn) mechanism (Chen et al., 2023a), which can be used to process long sequences by mitigating the high memory cost and slow processing time associated with standard self-attention in LLMs. S2-Attn divides input into sequence spans within self-attention modules and introduces shifted patterns for inter-span information exchange. This extends context length efficiently without extra computational costs and achieves near-baseline performance. It maintains the standard self-attention architecture during inference, ensuring compatibility with existing infrastructure and optimizations.
@@ 3.3 JMLR Architecture
However, Colbert matches queries and documents based on the retriever similarity, returning scores and documents. To enable the retriever system to return documents that are more helpful for providing accurate answers to Llama, we constructed the JMLR architecture. Let $\mathcal{Q}$ be a set of queries, $\mathcal{D}$ be a corpus of domain specific documents, and $\mathcal{A}$ be the set of options. Our task is to construct a function $f$ parameterized by $\theta$, which maps a question $Q\in\mathcal{Q}$ and a set of documents $D\subseteq\mathcal{D}$ to an answer $A\in\mathcal{A}$.
@@
where $F_{i}$ is calculated as:
@@
For JMLR model training, the dataset comprises QA pairs, with the input being a question and the output being the correct choice, including the reasoning part. The top 7 documents with the highest scores are aggregated and fed into the model during the training process. Prior to training, we extracted the top 30 documents based on scores from the retriever. During each iteration, 30 scores are generated by the retriever, and we perform weighted random sampling based on these scores, with higher scores having a greater probability of selection. This approach ensures that each document inputted into the LLM model is dynamic. The LLM will likely encounter different useful documents and some that are not useful. This step enhances the LLM’s robustness, helping it discern which knowledge from which documents is useful and which documents are irrelevant.
@@ 4.1 Dataset
We individually apply supervised finetuning on each QA dataset and subsequently assess their performance on the corresponding test sets, unless otherwise specified. MedQA and Amboss both offer detailed reasoning traces, such as extensive answers or explanations.
@@ 4.2 Baselines
For comparison with traditional domain pretraining method, we select several public available medical domain pretrained LLMs as baselines: (1) PMC-Llama, a Llama model further pretrained on medical research articles from PubMed Central (Wu et al., 2023); (2) Meditron, a LLM pretrained on clinical guidelines and research papers (Chen et al., 2023b).
To validate the benefit of joint training, we also compare JMLR with its naive version: RAG, which freezes the retriever and fetches the same document given a specific question.
@@ 4.3 Training Details
Both our training process and the conventional finetuning approach employ the AdamW optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\text{eps}=1\times 10^{-5}$. We implement a cosine learning rate schedule, incorporating a warmup phase that accounts for 10% of the training duration and decays the learning rate to 10% of its peak value. In alignment with the practices outlined in Llama 2-chat Touvron et al. (2023), our training employs a learning rate of $1\times 10^{-5}$, a weight decay factor of 0.1, and manages a batch size of 2. The finetuning phase spans 5 epochs for all iterations. However, we apply a distinct learning rate for ColBERT, set at $53e-$ 5. The optimization strategy for training ColBERT mirrors that used for Llama 2. Throughout these experiments, we utilize four A100 80 GB GPUs. The total training duration for our method amounts to 25 hours. Additionally, the pretraining phase on medical guidelines takes approximately 138 hours, while the finetuning phase is approximately 17 hours on 7B Llama. This comprehensive training approach ensures the effective adaptation and optimization of the model for specific medical applications.
@@ 5 Main Results
Model Parameter Open MedQA Amboss MMLU MedMCQA Average
Access
GPT-4 ? No 74.7 82.1 88.4 69.5 78.6
ChatGPT ? No 50.2 49.1 69.4 51.0 54.9
Meditron 70B Yes 60.7 76.4 73.6 65.1 68.9
RAG 13B Yes 59.9 76.9 69.9 64.2 67.7
JLMR 13B Yes 62.5 81.2 72.8 65.5 70.5
Table 2: Comparison between JMLR and SOTA open/closed access LLMs.
As shown in Table 2, our JLMR with 13B parameters outperforms previous SOTA open access model Meditron 70B parameters and even closed access model such as ChatGPT. In the following sections, we will explain with more detailed analysis.
@@ 5.1 Domain Retrieval > Domain Pretrain
Model MedQA Amboss MMLU MedMCQA Average
@@
RAG# 54.6 70.7 62.1 62.4 62.5
JLMR# 56.2 68.3 64.3 62.4 62.8
Table 3: Comparison between domain retrieval (#) and domain pretraining (*) on LLMs with 7B parameters.
Domain retrieval models (both RAG-7B and our JMLR-7B) outperform domain pretrained LLMs across all medical QA benchmarks, as shown in Table 3. In the MedQA and Amboss datasets, where JMLR-7B not only outperforms baselines but does so with a notable margin, scoring 56.2% in MedQA and 71.2% in Amboss. This is particularly significant when compared to Meditron-7B, which scores 47.9% and 50.1% in these datasets, respectively. The trend continues in MMLU-Medical and MedMcQA datasets, JMLR-7B achieved scores of 64.3% and 62.4%, respectively, surpassing the scores of PMC-Llama-7B (59.7% and 57.6%), and Meditron-7B (55.6% and 59.2%). Overall, the results illustrate that JMLR-7B model, on average, achieves about a 14% improvement in performance over its closest competitor, Meditron-7B. This highlights the effectiveness of our training method and the substantial advancements it brings to the field of medical benchmarking.
The enhanced performance of JMLR-7B can be partially attributed to its use of Llama-2 as the foundational model, showcasing significantly higher average performance than other pre-trained baselines. However, the distinct edge comes from our integrated training approach that combines retriever and LLM. This methodology not only further enhances Llama-2’s performance in medical benchmarks but also equips the retriever component with the ability to effectively source relevant and beneficial documents to aid the LLM in answering questions. In contrast, ChatGPT, when devoid of medical guidelines, has been known to make basic mistakes, such as recommending vaccinations to pregnant women—a practice that contradicts medical norms as documented in Table 5. Our model, with the support of guidelines, avoids such errors. This synergistic approach substantially reduces hallucinations and bolsters the model’s overall reliability.

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated a remarkable potential in medical knowledge acquisition and question-answering. However, LLMs can potentially hallucinate and yield factually incorrect outcomes, even with domain-specific pretraining. Previously, retrieval augmented generation (RAG) has limited success in addressing hallucinations. Unlike previous methods in RAG where the retrieval model was trained separately from the LLM, we introduce JMLR (for Jointly trains LLM and information Retrieval) during the fine-tuning phase. The synchronized training mechanism enhances JMLR’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions and reduces the demand for computational resources. We evaluated JMLR on the important medical question-answering application. Our experimental results demonstrate that JMLR-13B (70.5%) outperforms a previous state-of-the-art open-source model using conventional pre-training and fine-tuning Meditron-70B (68.9%) and Llama2-13B with RAG (67.7%) on a medical question-answering dataset. Comprehensive evaluations reveal JMLR-13B enhances reasoning quality and reduces hallucinations better than Claude3-Opus. Additionally, JMLR-13B (148 GPU hours) also trains much faster than Meditron-70B (42630 GPU hours). Through this work, we provide a new and efficient knowledge enhancement method for healthcare, demonstrating the potential of integrating retrieval and LLM training for medical question-answering systems.
 [^1] The code(https://github.com/believewhat/JMLR-Joint-Medical-LLM-and-Retrieval-Training), along with selected retrieval data that can be made public, is included in the supplementary material and will be made publicly accessible with CC-BY 4.0 license upon the paper’s acceptance.
@@ 1 Introduction
Figure 1: JMLR achieved the highest average accuracy across the MMLU-Medical, MedMcQA, MedQA, and Amboss datasets, utilizing only 148 GPU hours.
@@
Effective clinical decision-making relies on a logical diagnostic chain, which requires specialized knowledge that isn’t widely accessible, presenting a major healthcare challenge. Large Language Models (LLMs) show promise in making this critical medical knowledge more accessible (Peng et al., 2023; Yang et al., 2023). Traditionally, LLMs have been developed for general tasks using data from diverse online sources, leading to a lack of high-quality, domain-specific information, especially in medicine Wornow et al. (2023); Singhal et al. (2023a). Previous researches expand medical knowledge by continuing training general domain LLMs with domain-specific datasets. (Yunxiang et al., 2023; Zhang et al., 2023; Toma et al., 2023).
However, LLMs face the significant challenge of “hallucination”, where models generate plausible but incorrect or unverified information (Ji et al., 2023; Bang et al., 2023). Such errors raise serious concerns in healthcare, where accuracy is imperative (Ahmad et al., 2023). Hence, recent works proposed retrieval augmented generation (RAG): they first train a retriever to obtain relevant documents from a general domain corpus based on the input query and then train an LLM to generate a response based on the input query and the retrieved documents  (Lewis et al., 2020; Borgeaud et al., 2022a; Cheng et al., 2023; Hiesinger et al., 2023; Xiong et al., 2024). By retrieving domain-specific documents, RAG identifies relevant knowledge and provides contextual grounding for LLMs, thereby alleviating hallucination issues  (Shuster et al., 2021; Zhu et al., 2023).
Although RAG shows high accuracy in benchmarks for open-domain question answering, its effectiveness in specific domains is yet in challenge because retrievers trained in the general domain usually perform worse than those fine-tuned in the specific domain  (Gao et al., 2023; Zhao et al., 2022; Thakur et al., 2021). Fine-tuning retrievers requires document-query pairs, which may not be readily available for a specific domain. Labeling such medical pairs specifically for this purpose incurs additional time and financial costs. Moreover, RAG trains retriever and LLM separately. When training LLM with a frozen retriever, the retrieval may not be optimized to align with LLM to produce the correct answer (Asai et al., 2024; Yoran et al., 2023; Rubin et al., 2022). In this study, we introduce a novel method to jointly train retrieval and LLM. Although the method is generalizable, we focus on evaluating it in the important medical domain, where the accuracy of the generated content is crucial.
@@
To validate JMLR’s effectiveness in reducing computational resource requirements and fully utilizing given knowledge, we designed three experimental setups. These experiments evaluated whether direct fine-tuning, instead of traditional pretraining, would improve JMLR’s performance. We utilized external data resources for retrieval, including the MIMIC-IV dataset (Johnson et al., 2020), medical textbooks, and diverse medical documents for knowledge expansion. We experimented with different LLMs, including PMC-Llama, Meditron, GPT-3.5 with or without MedRAG. Our JMLR-7B has demonstrated superior results, achieving an impressive 62.3% accuracy, outperforming the traditional method’s 53.2% on the medical QA dataset. Furthermore, our 13B model (70.5%) surpasses the performance of both the open-source medical LLM (Meditron 70B: 68.9%) and closed-source general LLM (GPT-3.5: 54.9%). Moreover, evaluation by both GPT-4 and domain experts supports the superior performance of JMLR. We summarize our key contributions as follows:
  • We propose a novel method that jointly trains retrieval and LLMs, resulting in JMLR-13B surpassing the state-of-the-art 70B open-source model in several medical question-answering benchmarks.
  • Through comprehensive automated and human evaluations, we demonstrate that JMLR-13B consistently enhances models’ reasoning quality while significantly reducing hallucinations, achieving superior reasoning ability compared to Claude3-Opus.
  • JMLR shows promise to enhance domain-specific retrieval by fully eliminating additional human annotation.
@@ 3.1 Retriever
For information retrieval, we used the ColBERT model and the trained weight as the initial model. ColBERT utilizes BERT-based encoders to transform both queries and documents into bags of embeddings. A single BERT model is shared between the query and document encoders, but the inputs are distinguished by appending special tokens: $[Q]$ for queries and $[D]$ for documents. The text is tokenized for queries into BERT-based WordPiece tokens $q_{0},q_{1}\ldots q_{l}$, with the $[Q]$ token placed after BERT’s sequence-start token $[CLS]$. For documents, ColBERT segments a document d into its constituent tokens $d_{1},d_{2},\ldots d_{n}$, to which we prepend BERT’s start token $[CLS]$ followed by our special token $[D]$ that indicates a document sequence. Queries shorter than a predefined number of tokens $N_{q}$ are padded with BERT’s special $[mask]$ tokens to reach the length $N_{q}$, otherwise truncated. The padded sequence of input tokens is then passed through BERT. The resulting embeddings for queries ($E_{q}$) and documents ($E_{d}$) are computed as follows: $E_{q}$ is normalized after being processed through a CNN layer applied to BERT’s output, and $E_{d}$ undergoes a similar process but includes a filtering step to remove embeddings of punctuations.
@@ 3.2 LLM
We adapted Llama as our LLM model. The original Llama model supports only a maximum of 4k tokens, which poses a challenge when multiple documents are retrieved. To address this, we utilized the Shifted Sparse Attention (S2-Attn) mechanism (Chen et al., 2023a), which can be used to process long sequences by mitigating the high memory cost and slow processing time associated with standard self-attention in LLMs. S2-Attn divides input into sequence spans within self-attention modules and introduces shifted patterns for inter-span information exchange. This extends context length efficiently without extra computational costs and achieves near-baseline performance.
@@ 3.3 JMLR Architecture
However, Colbert matches queries and documents based on the retriever similarity, returning scores and documents. To enable the retriever system to return documents that are more helpful for providing accurate answers to Llama, we constructed the JMLR architecture. Let $\mathcal{Q}$ be a set of queries, $\mathcal{D}$ be a corpus of domain-specific documents, and $\mathcal{A}$ be the set of options. Our task is to construct a function $f$ parameterized by $\theta$, which maps a question $Q\in\mathcal{Q}$ and a set of documents $D\subseteq\mathcal{D}$ to an answer $A\in\mathcal{A}$.
@@
where M is the number of queries in the training set and $F_{i}$ is calculated as:
@@
For JMLR model training, the dataset comprises QA pairs, with the input being a question and the output being the correct choice, including the rationale part. The top 7 documents with the highest scores are aggregated and fed into the model during the training process. Prior to training, we extracted the top 30 documents based on scores from the retriever. During each iteration, 30 scores are generated by the retriever, and we perform weighted random sampling based on these scores, with higher scores having a greater probability of selection. This approach ensures that each document inputted into the LLM model is dynamic. The LLM will likely encounter different useful documents and some that are not useful. This step enhances the LLM’s robustness, helping it discern which knowledge from which documents is useful and which documents are irrelevant.
@@ 4.1 Dataset
We individually apply supervised finetuning on each QA dataset and subsequently assess their performance on the corresponding test sets, unless otherwise specified. MedQA and Amboss both offer detailed rationale with long explanations.
@@ 4.2 Baselines
For comparison with traditional domain pretraining method, we used Meditron, an open source LLM pretrained on clinical guidelines and research papers (Chen et al., 2023b). We also compared with other closed source LLMs such as OpenAI GPT-3.5, GPT-4 Achiam et al. (2023), and Anthropic-Claude3  Anthropic (2024) For the rationale evaluation experiment, we didn’t include Meditron because it cannot generate a valid rationale in our experiments. We didn’t include GPT-4 because it was used as the evaluator. To validate the benefit of joint training, we also compared JMLR with its naive version: RAG, which freezes the retriever and fetches the same document given a question.
4.3 Evaluation Metrics
For automated metrics, we used accuracy to evaluate the model’s ability to select the correct final choice. We evaluated the rationale quality using metrics UMLS-F and GPT-4 score. UMLS-F is a factuality metric based on F1 score between entities in gold explanation and generated explanation  [^2] More UMLS-F details can be found in Appendix A.4. GPT-4 score is a reference-free metric. GPT-4 was asked to rate the generated explanation on a Likert scale from 1 to 5 (higher better) along the 3 dimensions: question comprehension (e.g., indication the question has been understood), recall of knowledge (e.g., mention of a relevant and/or correct fact for answering the question), and medical reasoning (e.g., correct reasoning for answering the question) (Singhal et al., 2023a). We put GPT-4-score prompts in Table 8.
For human evaluation, we engaged three medical professionals to review and assess a small sample of 20 questions drawn from the test sets. These doctors were asked to provide their preference for the rationales generated by two models (GPT-3.5 and JMLR-13b) given the gold rationale.
To validate the reliability of the selected automated metrics, we calculated the Cohen’s Kappa between UMLS-F and expert preference, as well as Kappa between the GPT-4 score and expert preference, obtaining results of 0.69 and 0.81, respectively. These results indicate that both metrics exhibit substantial to almost perfect agreement with human evaluation.
@@ 4.4 Training Details
Both our training process and the conventional finetuning approach employ the AdamW optimizer, with $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\text{eps}=1\times 10^{-5}$. We implement a cosine learning rate schedule, incorporating a warmup phase that accounts for 10% of the training duration and decays the learning rate to 10% of its peak value. In alignment with the practices outlined in Llama 2-chat Touvron et al. (2023), our training employs a learning rate of $1\times 10^{-5}$, a weight decay factor of 0.1, and manages a batch size of 2. The finetuning phase spans 5 epochs for all iterations. However, we apply a distinct learning rate for ColBERT, set at $53e-$ 5. The optimization strategy for training ColBERT mirrors that used for Llama 2. Throughout these experiments, we utilize four A100 80 GB GPUs.
@@ 5 Main Results
As shown in Table 2, our JMLR with 13B parameters outperforms previous SOTA open access model Meditron 70B parameters and even closed access model such as GPT-3.5.
Model MedQA Amboss MMLU-M MedMCQA AVG
GPT-3.5 50.2 49.1 69.4 51.0 54.9
GPT-4 74.7 82.1 88.4 69.5 78.6
Meditron-70B 60.7 76.4 73.6 65.1 68.9
RAG-13B 59.9 76.9 69.9 64.2 67.7
JMLR-13B 62.5 81.2 72.8 65.5 70.5
Table 2: Comparison between JMLR and SOTA open (Meditron, RAG, JMLR) or closed (GPT-3.5/4) LLMs.
@@ 5.1 Domain Retrieval > Domain Pretrain
Domain retrieval models (both RAG-7B and our JMLR-7B) outperform domain-pretrained LLMs across all medical QA benchmarks, as shown in Table 3. In the MedQA and Amboss datasets, where JMLR-7B not only outperforms baselines but does so with a notable margin, scoring 51.3% in MedQA and 68.3% in Amboss. This is particularly significant when compared to Meditron-7B, which scores 47.9% and 50.1% in these datasets, respectively. The trend continues in MMLU-Medical and MedMcQA datasets, JMLR-7B achieved scores of 65.3% and 64.1%, surpassing the scores of Meditron-7B (55.6% and 59.2%). Overall, the results illustrate that JMLR-7B model, on average, achieves about a 17% improvement in performance over its closest competitor, Meditron-7B. This highlights the effectiveness of our training method and the substantial advancements it brings to the field of medical benchmarking.
The enhanced performance of JMLR-7B can be partially attributed to its use of Llama-2 as the foundational model, showcasing significantly higher average performance than other pre-trained baselines. However, the distinct edge comes from our integrated training approach that combines retriever and LLM. This methodology not only further enhances Llama-2’s performance in medical benchmarks but also equips the retriever component with the ability to effectively source relevant and beneficial documents to aid the LLM in answering questions. In contrast, GPT-3.5, when devoid of medical guidelines, has been known to make basic mistakes, such as recommending vaccinations to pregnant women—a practice that contradicts medical norms as documented in Table 6. Our model, with the support of guidelines, avoids such errors. This synergistic approach substantially reduces hallucinations and bolsters the model’s overall reliability.
Model MedQA Amboss MMLU-M MedMCQA AVG
@@
RAG# 47.3 50.7 63.8 62.1 55.9
JMLR# 51.3 68.3 65.3 64.1 62.3
Table 3: Comparison between domain retrieval (#) and domain pretraining (*) using models of the same 7B size and domain data (Open Guidelines).